# FBSDiff: Plug-and-Play Frequency Band Substitution of Diffusion Features for Highly Controllable Text-Driven Image Translation

### v1

@@ Abstract
Large-scale text-to-image diffusion models have been a revolutionary milestone in the evolution of generative AI and multimodal technology, allowing extraordinary image generation based on natural-language text prompts. However, the issue of lacking controllability of such models restricts their practical applicability for real-life content creation, for which attention has been focused on leveraging a reference image to control text-to-image synthesis. Due to the close correlation between the reference image and the generated image, this problem can also be regarded as the task of manipulating (or editing) the reference image as per the text, namely text-driven image-to-image translation. This paper contributes a novel, concise, and efficient approach that adapts the pre-trained large-scale text-to-image (T2I) diffusion model to the image-to-image (I2I) paradigm in a plug-and-play manner, realizing high-quality and versatile text-driven I2I translation without any model training, model fine-tuning, or online optimization process. To guide T2I generation with a reference image, we propose to model diverse guiding factors with correspondingly different frequency bands of diffusion features in the DCT spectral space, and accordingly devise a novel frequency band substitution layer that dynamically substitutes a certain DCT frequency band of the diffusion features with the corresponding counterpart of the reference image along the reverse sampling process. We demonstrate that our method flexibly enables highly controllable text-driven I2I translation both in the guiding factor and guiding intensity of the reference image, simply by tuning the type and bandwidth of the substituted frequency band, respectively. Extensive qualitative and quantitative experiments verify the superiority of our approach over related methods in I2I translation visual quality, versatility, and controllability.
@@
Figure 1: Based on the pre-trained text-to-image diffusion model, FBSDiff enables efficient text-driven image-to-image translation by proposing a plug-and-play reference image guidance mechanism. It allows flexible control over different guiding factors (e.g., image appearance, image layout, image contours) of the reference image to the T2I generation result, simply by dynamically substituting different types of DCT frequency bands during the reverse sampling process of the diffusion model. Better viewed with zoom-in.
@@ I Introduction
As a typical application of the booming multimodal technology, text-driven I2I translation is an appealing computer vision problem that aims to translate a source (reference) image as per a text prompt. It can also be treated as a more controllable form of text-to-image (T2I) synthesis: controlling the T2I generation with a reference image. Since the advent of CLIP [1] bridging vision and language with large-scale contrastive pre-training, attempts have been made to instruct image manipulation with text by combining CLIP with generative models. VQGAN-CLIP [2] pioneers text-driven image translation by optimizing VQGAN [3] latent image embedding using the CLIP text-image similarity loss. DiffusionCLIP [4] fine-tunes diffusion model [5] under the CLIP loss to manipulate an image as per a text. DiffuseIT [6] combines VIT-based structure loss [7] and CLIP-based semantic loss to guide the diffusion model’s reverse sampling process via manifold constrained gradient [8], synthesizing translated image that complies with the target text while maintaining the structure of the reference image. However, these methods are not competitive in visual quality of the generation results, due to the limited model capacity of the backbone generative model as well as unstability caused by the online fine-tuning process.
To promote image translation visual quality, efforts have been made to train large models on massive data. InstructPix2Pix [9] employs GPT-3 [10] and Stable Diffusion [11] to synthesize huge amounts of paired training data, based on which trains a supervised text-driven I2I mapping for general image manipulation task. Design Booster [12] trains a latent diffusion model [11] conditioned on a joint representation that fuses both text embedding and image embedding, realizing layout-preserved text-driven I2I translation. Nevertheless, these methods are computationally intensive due to the need for training large models on immense data.
@@
The former type of fine-tuning-based methods represented by SINE [13] and Imagic [14] fine-tune the pre-trained T2I diffusion model to reconstruct the reference image before manipulating it with a target text. These methods require separate fine-tuning of the entire diffusion model for each time of image editing, which is less efficient and prone to underfitting or overfitting to the reference image.
@@
The latter type of inversion-based methods invert the reference image into the diffusion model’s Gaussian noise space and then generate the translated image via the reverse sampling process guided by the target text. A pivotal challenge of this pipeline is that the sampling trajectory may severely deviate from the inversion trajectory due to the error accumulation caused by the classifier-free guidance technique [15], which impairs the correlation between the reference and the translated image. To remedy this issue, Null-text Inversion [16] optimizes the unconditional null-text embedding to calibrate the sampling trajectory step by step. Prompt Tuning Inversion [17] proposes to minimize trajectory divergence by learning to encode the reference image into a learnable prompt embedding. Similarly, StyleDiffusion [18] opts to optimize the “value” embedding of the cross-attention layer to encode the visual information of the reference image. Pix2Pix-zero [19] penalizes trajectory deviation by matching cross-attention maps between the two trajectories with least-square loss. These methods apply per-step online optimization to calibrate the whole sampling trajectory, introducing additional computational cost and time overhead. Moreover, most of these methods rely on the cross-attention control technique introduced by Prompt-to-Prompt [20] for image structure preservation. This makes them rely on a paired source text of the reference image for word-to-word alignment, which is not flexible or even available in most cases. Plug-and-Play (PAP) [21] proposes to use feature maps and self-attention maps extracted from the denoising U-Net internal layers to maintain image structure, realizing optimization-free text-driven I2I translation. However, the algorithm is sensitive to specific layers for feature extraction and feature injection, and the feature extraction process is also time-consuming.
@@
In this paper, we propose a concise and efficient approach termed FBSDiff, realizing plug-and-play and highly controllable text-driven I2I translation from a frequency-domain perspective. To guide T2I generation with a reference image, a key missing ingredient of existing methods is the mechanism to control the guiding factor (e.g., image appearance, image layout, and image contours) and guiding intensity of the reference image. Since different guiding factors are difficult to isolate in the spatial domain, we consider decomposing them in the frequency domain by modeling different guiding factors with different frequency bands of diffusion features in the Discrete Cosine Transform (DCT) spectral space. Based on this motivation, we propose an inversion-based text-driven I2I translation framework featured with a novel frequency band substitution mechanism, realizing plug-and-play and highly controllable reference image guidance by dynamically substituting a certain DCT frequency band of diffusion features with the corresponding counterpart of the reference image along the reverse sampling process. As displayed in Fig. 1, T2I generation with appearance and layout control, pure layout control, and contour control of the reference image can be respectively realized by transplanting low-frequency band, mid-frequency band, and high-frequency band between diffusion features, enabling controllable and versatile text-driven I2I translation.
@@
The strengths of our method are fourfold: (I) dynamic reference image control: it allows plug-and-play text-driven I2I translation at inference time; (II) conciseness and efficiency: our method dispenses with the need for the paired source text of the reference image as well as cumbersome attention modulation process as compared with existing advanced methods, all while achieving leading I2I translation performance; (III) more methodologically generic : our method transplants frequency band of the denoised features along the reverse sampling trajectory, requiring no access to any internal features of the denoising network, and thus decouples with the specific diffusion model backbone architecture as contrasted with existing methods; (IV) high controllability: our method allows flexible control over the guiding factor and guiding intensity of the reference image simply by tuning the type and bandwidth of the substituted frequency band. The effectiveness of our method is fully demonstrated both qualitatively and quantitatively. To summarize, we make the following key contributions:
@@
  • We propose a novel frequency band substitution technique, realizing plug-and-play text-driven I2I translation without any model training, fine-tuning, or online optimization process.
@@ II-A Diffusion Model
Since the advent of DDPM [5], diffusion model has soon dominated the family of generative models [22]. Afterwards, much progress have been made to improve diffusion model in both methodology and application. DDIM [23] and its variants [24, 24] substantially accelerate diffusion model sampling process to tens of times with only marginal drop in generation quality, promoting its practicability dramatically. Palette [25] extends diffusion model from unconditional image generation to the realm of conditional image synthesis, opening the door of diffusion-based image-to-image translation. With the advancement of multimodal technology, large-scale T2I diffusion models [26, 27, 28] are proposed to generate high-resolution images with open-domain text prompts, bringing content creation to an unprecedented level. To lower computational overhead of large-scale T2I model, Latent Diffusion Model (LDM) [11] proposes to transfer diffusion model from high-dimension pixel space to low-dimensional feature space, contributing the most widely used architecture in AIGC industry. To introduce more controllability to T2I synthesis, ControlNet [29] and T2i-adapter [30] add spatial control to T2I diffusion models by training a control module of the denoising U-Net conditioned on certain image priors (e.g., canny edges, depth maps, human key points, etc.). SDXL [31] and DiTs [32] propose Transformer [33] based backbone denoising network, improving T2I diffusion model to larger capacity. Up to now, diffusion model has been applied to a wide variety of vision fields such as image super-resolution [34], image inpainting [35], image colorization [36], semantic segmentation [37], point cloud generation [38], video synthesis [39], 3D reconstruction [40], etc, and is still making rapid progress in theory and potential applications.
@@
Figure 2: Overview of FBSDiff. Based on the pre-trained latent diffusion model (LDM), FBSDiff starts with an inversion process that inverts the reference image into the Gaussian noise space of the LDM, based on which a reconstruction process is applied to reconstruct the reference image from the inverted Gaussian noise, providing intermediate denoising results as pivotal guidance features. The guidance features are used to guide the text-driven sampling process of the LDM to exert reference image control onto the T2I result, which is realized by dynamically transplanting certain DCT frequency bands from the reconstruction feature trajectory into the sampling feature trajectory. The dynamic DCT frequency band transplantation is realized by our proposed frequency band substitution layer (FBS layer) in a plug-and-play manner.
@@
Figure 3: Illustration of the proposed frequency band substitution (FBS) layer. The FBS layer takes in two diffusion features and substitutes a certain DCT frequency band of one feature with the corresponding frequency band of the other feature. This is realized by converting the two features into the frequency domain via 2D DCT, extracting and transplanting a certain DCT frequency band, and converting the manipulated DCT features back to spatial domain via 2D IDCT. The frequency band extraction and transplantation process are implemented with binary masking.
@@ II-B Computer Vision in Frequency Perspective
Neural network models are mostly applied to tackle vision tasks in spatial or temporal domain, some research reveals that model performance can also be boosted from frequency domain. For example, Ghosh et al. [41] introduce DCT to convolutional neural network for image classification, accelerating network convergence speed. Xie et al. [42] propose a frequency-aware dynamic network for lightweight image super-resolution. Cai et al. [43] impose Fourier frequency spectrum consistency to image translation tasks, achieving better identity preservation ability. FreeU [44] improves T2I generation quality by selectively enhancing or depressing different frequency components of diffusion features inside the denoising U-Net model. ILVR [45] proposes to fuse the low-frequency information of the forward diffusion process into the reverse sampling process for conditioned image synthesis. Our method differs with ILVR in that ILVR simulates low-pass filtering with simple feature downsampling and upsampling and performs information fusion in the spatial domain, while our method explicitly extracts and transplants frequency bands of diffusion features in the DCT domain. FCDiffusion [46] shares similar spirit of frequency-based control of T2I diffusion model with our method. However, FCDiffusion relies on training multiple frequency control branches to realize versatile control effects, while our method achieves versatility and high controllability in both guiding factor and guiding intensity of the reference image in a training-free and plug-and-play manner.
@@ III-A Overall Architecture
Established on the pre-trained Latent Diffusion Model (LDM), FBSDiff adapts it from T2I generation to the realm of text-driven I2I translation with our proposed plug-and-play reference image guidance mechanism: dynamic frequency band substitution, which efficiently and flexibly realizes controllable guiding factor and guiding intensity of the reference image to the T2I generation result.
@@
Meanwhile, an equal-length sampling trajectory is applied in parallel with the reconstruction trajectory for T2I synthesis. The sampling trajectory is also a $T$-step DDIM sampling process that progressively denoises a randomly initialized Gaussian noise feature $\tilde{x}_{T}\sim\mathcal{N}(0,I)$ into $\tilde{x}_{0}$ conditioned on the text embedding $v$ of the target text prompt. To amplify the effect of text guidance, we employ classifier-free guidance technique [15] by interpolating conditional (target text) and unconditional (null text) noise prediction at each time step with a guidance scale $\omega$ during the sampling trajectory:
@@
Due to the inherent property of DDIM inversion and DDIM sampling, the reconstruction trajectory forms a deterministic denoising mapping towards the reference image, where its intermediate denoising results $\{\hat{z}_{t}\}$ can be used as pivotal guidance features to calibrate the corresponding counterpart $\{\tilde{z}_{t}\}$ along the sampling trajectory. Therefore, the correlation between the reference image and the generated image can be established, and thus realizing text-driven I2I translation. Specifically, we implement feature calibration by inserting a plug-and-play frequency band substitution (FBS) layer in between the reconstruction trajectory and the sampling trajectory. FBS layer substitutes a certain frequency band of $\tilde{z}_{t}$ in the sampling trajectory with the corresponding frequency band of $\hat{z}_{t}$ in the reconstruction trajectory along the reverse sampling process. Such frequency band substitution effectively and efficiently imposes guidance of the reference image to the T2I synthesis process. Both the guiding factor (e.g., image appearance, image layout, image contours) and guiding intensity can be flexibly controllable simply by tuning the type and bandwidth of the substituted frequency band, respectively.
To improve I2I translation visual quality, we partition the sampling trajectory into a calibration phase and a non-calibration phase, separated by the time step $\lambda T$. In the former calibration phase ($\tilde{z}_{T}\rightarrow\tilde{z}_{\lambda T}$), dynamic frequency band substitution is applied at each time step for smooth calibration of the sampling process; in the latter non-calibration phase ($\tilde{z}_{\lambda T-1}\rightarrow\tilde{z}_{0}$), we remove FBS layer to avoid over-constrained sampling result, fully unleashing the generative power of the pre-trained T2I model to improve image generation quality. Here $\lambda$ denotes the ratio of the length of the non-calibration phase to that of the entire sampling trajectory.

### v2

@@ Abstract
Large-scale text-to-image diffusion models have been a revolutionary milestone in the evolution of generative AI and multimodal technology, allowing wonderful image generation with natural-language text prompt. However, the issue of lacking controllability of such models restricts their practical applicability for real-life content creation. Thus, attention has been focused on leveraging a reference image to control text-to-image synthesis, which is also regarded as manipulating (or editing) a reference image as per a text prompt, namely, text-driven image-to-image translation. This paper contributes a novel, concise, and efficient approach that adapts pre-trained large-scale text-to-image (T2I) diffusion model to the image-to-image (I2I) paradigm in a plug-and-play manner, realizing high-quality and versatile text-driven I2I translation without any model training, model fine-tuning, or online optimization process. To guide T2I generation with a reference image, we propose to decompose diverse guiding factors with different frequency bands of diffusion features in the DCT spectral space, and accordingly devise a novel frequency band substitution layer which realizes dynamic control of the reference image to the T2I generation result in a plug-and-play manner. We demonstrate that our method allows flexible control over both guiding factor and guiding intensity of the reference image simply by tuning the type and bandwidth of the substituted frequency band, respectively. Extensive qualitative and quantitative experiments verify superiority of our approach over related methods in I2I translation visual quality, versatility, and controllability. The code is publicly available at: https://github.com/XiangGao1102/FBSDiff.
@@
Figure 1: Based on the pre-trained text-to-image diffusion model, FBSDiff enables efficient text-driven image-to-image translation by proposing a plug-and-play reference image guidance mechanism. It allows flexible control over different guiding factors (e.g., image appearance, image layout, image contours) of the reference image to the T2I generated image, simply by dynamically substituting different types of DCT frequency bands during the reverse sampling process of the diffusion model. Better viewed with zoom-in.
@@ I Introduction
As a typical application of the booming multimodal technology, text-driven I2I translation is an appealing computer vision problem that aims to translate a reference image as per a text prompt. It extends text-to-image (T2I) synthesis to more controllability by controlling T2I generation result with a reference image. Since the advent of CLIP [1] bridging vision and language through large-scale contrastive pre-training, attempts have been made to instruct image manipulation with text by combining CLIP with generative models. VQGAN-CLIP [2] pioneers text-driven image translation by optimizing VQGAN [3] latent image embedding with CLIP text-image similarity loss. DiffusionCLIP [4] fine-tunes diffusion model [5] under the same CLIP loss to manipulate an image as per a text. DiffuseIT [6] combines VIT-based structure loss [7] and CLIP-based semantic loss to guide diffusion model’s reverse sampling process via manifold constrained gradient [8], synthesizing translated image that complies with the target text while maintaining the structure of the reference image. However, these methods are not competitive in generation visual quality due to the limited model capacity of backbone generative model as well as the inherent unstability caused by online fine-tuning or optimization process.
To promote image translation visual quality, efforts have been made to train large models on massive data. InstructPix2Pix [9] employs GPT-3 [10] and Stable Diffusion [11] to synthesize huge amounts of paired training data, based on which trains a supervised text-driven I2I mapping for general image manipulation task. Design Booster [12] trains a latent diffusion model [11] conditioned on both text embedding and image embedding, realizing layout-preserved text-driven I2I translation. Nevertheless, these methods are computationally intensive in training large models from scratch and less efficient in collecting immense training data.
@@
The former type of fine-tuning-based methods represented by SINE [13] and Imagic [14] fine-tune the pre-trained T2I diffusion model to reconstruct an input reference image before manipulating it with a target text. These methods require separate fine-tuning of the entire diffusion model for each time of image manipulation, which is less efficient and prone to underfitting or overfitting to the reference image.
@@
The latter type of inversion-based methods invert reference image into diffusion model’s Gaussian noise space and then generate the translated image via the reverse sampling process guided by the target text. A pivotal challenge of this pipeline is that the sampling trajectory may severely deviate from the inversion trajectory due to the error accumulation caused by the classifier-free guidance technique [15], which severely impairs the correlation between the reference image and the translated image. To remedy this issue, Null-text Inversion [16] optimizes the unconditional null-text embedding to calibrate the sampling trajectory step by step. Prompt Tuning Inversion [17] proposes to minimize trajectory divergence with an optimization to encode the reference image into a learnable prompt embedding. Similarly, StyleDiffusion [18] opts to optimize the “value” embedding of the cross-attention layer as the visual encoding of the reference image. Pix2Pix-zero [19] penalizes trajectory deviation by matching cross-attention maps between the two trajectories with least-square loss. These methods apply per-step online optimization to calibrate the whole sampling trajectory, introducing additional computational cost and time overhead. Moreover, most of these methods adopt the cross-attention control technique introduced in Prompt-to-Prompt [20] for image structure preservation. This makes them rely on a paired source text of the reference image, which is not flexible or even available in most cases. Plug-and-Play (PAP) [21] proposes to leverage feature maps and self-attention maps extracted from internal layers of the denoising U-Net to maintain image structure, realizing optimization-free text-driven I2I translation. However, the algorithm is sensitive to specific layer selection, and the feature extraction process is also time-consuming.
@@
In this paper, we propose a concise and efficient approach termed FBSDiff, realizing plug-and-play and highly controllable text-driven I2I translation from a frequency-domain perspective. To guide T2I generation with a reference image, a key missing ingredient of existing methods is the mechanism to control the guiding factor (e.g., image appearance, layout, contours) and guiding intensity of the reference image. Since different image guiding factors are difficult to isolate in the spatial domain, we consider decomposing them in the frequency domain by modeling them with different frequency bands of diffusion features in the Discrete Cosine Transform (DCT) spectral space. Based on this motivation, we propose an inversion-based text-driven I2I translation framework featured with a novel frequency band substitution mechanism, which efficiently enables reference image guidance of the T2I generation by dynamically substituting a certain DCT frequency band of diffusion features with the corresponding counterpart of the reference image along the reverse sampling process. As displayed in Fig. 1, T2I generation with appearance and layout control, pure layout control, and contour control of the reference image can be respectively realized by transplanting low-frequency band, mid-frequency band, and high-frequency band between diffusion features, allowing versatile and highly controllable text-driven I2I translation.
@@
The strengths of our method are fourfold: (I) plug-and-play efficiency: our method extends pre-trained T2I diffusion model to the realm of I2I in a plug-and-play manner; (II) conciseness: our method dispenses with the need for the paired source text of the reference image as well as cumbersome attention modulation process as compared with existing advanced methods, all while achieving leading I2I translation performance; (III) model generalizability: our method transplants frequency band of diffusion features along the reverse sampling trajectory, requiring no access to any internal features of the denoising network, and thus decouples with the specific diffusion model backbone architecture as compared with existing methods; (IV) controllability: our method allows flexible control over the guiding factor and guiding intensity of the reference image simply by tuning the type and bandwidth of the substituted frequency band. The effectiveness of our method is fully demonstrated both qualitatively and quantitatively. To summarize, we make the following key contributions:
@@
  • We propose a novel frequency band substitution technique, realizing plug-and-play text-driven I2I translation without any model training, model fine-tuning, and online optimization process.
@@ II-A Diffusion Model
Since the advent of DDPM [5], diffusion model has soon dominated the family of generative models [22]. Afterwards, much progress have been made to improve diffusion model in both methodology and application. DDIM [23] and its variants [24, 24] accelerate diffusion model sampling process to tens of times with only marginal drop in generation quality, promoting its practicability dramatically. Palette [25] extends diffusion model from unconditional image generation to the realm of conditional image synthesis, opening the door of diffusion-based image-to-image translation. With the advancement of multimodal technology, large-scale T2I diffusion models [26, 27, 28] are proposed to generate high-resolution images with open-domain text prompts, bringing content creation to an unprecedented level. To lower computational overhead of large-scale T2I model, Latent Diffusion Model (LDM) [11] proposes to transfer diffusion model from high-dimension pixel space to low-dimensional feature space, contributing the most widely used architecture in AIGC industry. To introduce more controllability to T2I synthesis, ControlNet [29] and T2i-adapter [30] add spatial control to T2I diffusion models by training a control module of the denoising U-Net conditioned on certain image priors (e.g., canny edges, depth maps, human key points, etc.). SDXL [31] and DiTs [32] propose Transformer [33] based backbone denoising network, improving T2I diffusion model to larger capacity. Up to now, diffusion model has been applied to a wide variety of vision fields such as image super-resolution [34], image inpainting [35], image colorization [36], semantic segmentation [37], point cloud generation [38], video synthesis [39], 3D reconstruction [40], etc, and is still making rapid progress in theory and potential applications.
@@
Figure 2: Overview of FBSDiff. Based on the pre-trained latent diffusion model (LDM), FBSDiff starts with an inversion trajectory that inverts reference image into the LDM Gaussian noise space, then a reconstruction trajectory is applied to reconstruct the reference image from the inverted Gaussian noise, providing intermediate denoising results as pivotal guidance features. The guidance features are leveraged to guide the text-driven sampling trajectory of the LDM to exert reference image control, which is realized by dynamically transplanting certain DCT frequency bands from diffusion features along the reconstruction trajectory into the corresponding features along the sampling trajectory. The dynamic DCT frequency band transplantation is implemented in a plug-and-play manner with our proposed frequency band substitution layer (FBS layer).
@@
Figure 3: Illustration of the proposed frequency band substitution (FBS) layer. The FBS layer takes in two diffusion features and substitutes a certain frequency band of one feature with the corresponding frequency band of the other feature. This is realized by converting the two diffusion features into the frequency domain via 2D DCT, extracting and transplanting a certain DCT frequency band, and converting the fused DCT features back to spatial domain via 2D IDCT. The frequency band extraction and transplantation are implemented with binary masking.
@@ II-B Computer Vision in Frequency Perspective
Deep neural network models are mostly applied to tackle vision tasks in spatial or temporal domain, some research reveals that model performance can also be boosted from frequency domain. For example, Ghosh et al. [41] introduce DCT to convolutional neural network for image classification, accelerating network convergence speed. Xie et al. [42] propose a frequency-aware dynamic network for lightweight image super-resolution. Cai et al. [43] impose Fourier frequency spectrum consistency to image translation tasks, achieving better identity preservation ability. FreeU [44] improves T2I generation quality by selectively enhancing or depressing different frequency components of diffusion features inside the denoising U-Net model. ILVR [45] proposes to fuse low-frequency information of the forward diffusion process into the reverse sampling process for conditioned image synthesis. Our method differs with ILVR in that ILVR simulates low-pass filtering with simple feature downsampling and upsampling and performs information fusion in the spatial domain, while our method explicitly extracts and transplants frequency bands of diffusion features in the DCT domain. FCDiffusion [46] shares similar spirit of frequency-based control of T2I diffusion model with our method. However, FCDiffusion relies on training multiple frequency control branches to realize versatile control effects, while our method achieves versatility and high controllability in both guiding factor and guiding intensity of the reference image in a training-free and plug-and-play manner.
@@ III-A Overall Architecture
Established on the pre-trained Latent Diffusion Model (LDM), FBSDiff adapts it from T2I generation to the realm of text-driven I2I translation with our proposed plug-and-play reference image guidance mechanism: dynamic frequency band substitution, which efficiently realizes flexible control over both guiding factor and guiding intensity of the reference image to the T2I generated image.
@@
Meanwhile, an equal-length sampling trajectory is applied in parallel with the reconstruction trajectory for T2I synthesis. The sampling trajectory is also a $T$-step DDIM sampling process that progressively denoises a randomly initialized Gaussian noise feature $\tilde{x}_{T}\sim\mathcal{N}(0,I)$ into $\tilde{x}_{0}$ conditioned on the text embedding $v$ of the target text prompt. To amplify the effect of text guidance, we employ classifier-free guidance technique [15] by interpolating conditional (target text) and unconditional (null text) noise prediction at each time step with a guidance scale $\omega$ along the sampling process:
@@
Due to the inherent property of DDIM inversion and DDIM sampling, the reconstruction trajectory forms a deterministic denoising mapping towards the reference image, during which the intermediate denoising results $\{\hat{z}_{t}\}$ can function as pivotal guidance features to calibrate the corresponding counterparts $\{\tilde{z}_{t}\}$ along the sampling trajectory. Thus, correlation between the reference image and the generated image can be established to allow for text-driven I2I translation. Specifically, we implement feature calibration by inserting a plug-and-play frequency band substitution (FBS) layer in between the reconstruction trajectory and the sampling trajectory. FBS layer substitutes a certain frequency band of $\tilde{z}_{t}$ in the sampling trajectory with the corresponding frequency band of $\hat{z}_{t}$ in the reconstruction trajectory along the reverse sampling process. The frequency band substitution effectively and efficiently imposes guidance of the reference image to the T2I synthesis process. Both the guiding factor (e.g., image appearance, image layout, image contours) and guiding intensity of the reference image can be flexibly controlled simply by tuning the type and bandwidth of the substituted frequency band, respectively.
To improve I2I translation visual quality, we partition the sampling process into a calibration phase and a non-calibration phase, separated by the time step $\lambda T$. In the former calibration phase ($\tilde{z}_{T}\rightarrow\tilde{z}_{\lambda T}$), dynamic frequency band substitution is applied at each time step for smooth calibration of the sampling trajectory; in the latter non-calibration phase ($\tilde{z}_{\lambda T-1}\rightarrow\tilde{z}_{0}$), we remove FBS layer to avoid over-constrained sampling result, fully unleashing the generative power of the pre-trained T2I model to improve image generation quality. Here $\lambda$ denotes the ratio of the length of the non-calibration phase to that of the entire sampling trajectory.